OpenStudy (calculusxy):

MEDAL for checking an answer with Scientific Notations operation! Question posted below...

OpenStudy (calculusxy):

\[\huge 5.3 \times 10^{-5} \times 8.8 \times 10^{-3}\]

OpenStudy (calculusxy):

\[\huge Answer \] \[\large 4.664 \times 10^{-7}\]

hero (hero):

Result can only be accurate to two decimal places

hero (hero):

Also you added exponents incorrectly

OpenStudy (calculusxy):

I don't understand ...

hero (hero):

When you multiply using decimals you have to count the total number of digits that occur after the decimal point. In this case there are two . One for each number . So your answer can only be accurate to two decimal places.

OpenStudy (calculusxy):

Can you explain this to me more with the question? I can't grasp what you're trying to say

hero (hero):

One moment

OpenStudy (calculusxy):

@Hero

OpenStudy (calculusxy):

so where was my mistake?

hero (hero):

You have three digits expressed after the decimal in your answer

hero (hero):

µo you have to use rules of exponents When multiplying powers.

OpenStudy (calculusxy):

but the answer after i multiplied was 46.64.. so i moved it one space to the left to make it between 1 and 10 and so i got 4.664.

hero (hero):

The exponent part is definitely incorrect though

OpenStudy (calculusxy):

can you explain to me how?

hero (hero):

Because you add exponents when multiplying powers

OpenStudy (calculusxy):

i did .

OpenStudy (calculusxy):

didn't i?

hero (hero):

You know, I realise now that your result IS correct

OpenStudy (calculusxy):

okay thank you. can you make sure another answer of mine? \[\large 6.4 \times 10^{-1} \times 6.6 \times 10^{-9}\]

OpenStudy (calculusxy):

would the answer be \[\large 4.224 \times 10^{-9}\]

OpenStudy (calculusxy):

@Hero

hero (hero):

If you followed the same format as you did with your previous problem, it should be correct. You can check with wolfram

OpenStudy (calculusxy):

Okay thank you so much!!!!!!